Mrs. Mack teaches three art classes, each having 30 students. She has a supply of 150 sheets of special paper for an upcoming project. For the project, each student in each class will use one sheet of paper.
What fraction of the paper supply will be used by the three classes? The fraction in lowest terms.

Answer by josmiceli(19441)

90/150 = 9/15
9/15 = 3/5
3/5 of the paper supply is used